# Teaching an Old Dog New Tricks: Verifiable FHE Using Commodity Hardware

### v1

@@ 9.7. Analysis of the Remaining Attack Surface
Argos primarily defends against microarchitectural side channels, which constitute the vast majority of TEE attacks so far (43 published according to (Li et al., 2024)). While Argos is vulnerable to fault injection (Qiu et al., 2019; Murdock et al., 2020; Kenjar et al., 2020) and Rowhammer attacks (less studied on TEEs), these threats are not fatal – software can be hardened against fault injections (Breier and Hou, 2022), and no Rowhammer attacks have demonstrated the capabilities needed to compromise Argos (gaining hypervisor privilege on a hardware VM). Regarding physical attacks, Argos protects against ColdBoot (Halderman et al., 2009) and physical side channels on the CPU (Lee et al., 2020a), but remains vulnerable to physical fault injection attacks (Chen et al., 2021; Buhren et al., 2021). The TPM only exposes a minimal attack surface for physical side channels, and modern TPMs are now directly integrated into the SoC, significantly raising the bar for these attacks (Jacob et al., 2023). Finally, our TCB might still contain bugs, and attacks that exploit implementation mistakes have been demonstrated on the Intel DRoT and the associated secure coprocessor (Wojtczuk and Rutkowska, 2009; Wojtczuk et al., 2009; Ermolov and Goryachy, 2017; Corporation, 2019).
@@
SC: Side Channel, P: Protected, V: Vulnerable
@@ 10. Related Work
TPMs and Hypervisor-based TEEs. Trusted platform modules (TPMs) combined with dynamic root-of-trust technologies have long served as dedicated hardware to enforce platform integrity. Flicker (McCune et al., 2008)) proposes to use these two technologies combined to isolate and attest small pieces of application logic, or PALs while reducing the software TCB to a handful of lines of code. No keys ever leave the TPM in Flicker, which gives it the same resistance to microarchitectural side channels as Argos (even if not discussed in the paper at the time). However, their complete reliance on hardware mechanisms creates significant overheads between 2-3 orders of magnitude. Virtualization technologies (making their appearance in processors around 2008) aim at isolating different trusted domains from each other in the form of virtual machines (VMs). They create a new privilege execution mode dedicated to the hypervisor and provide hardware primitives such as extended (or nested) page tables to enforce memory isolation between different security domains. These technologies are quickly identified as an opportunity to virtualize TPM (Perez et al., 2006) and provide confidentiality and integrity for trusted domains with minimal performance overhead. TrustVisor (McCune et al., 2010) is the first work to introduce this hypervisor-based architecture which is eventually adopted by cloud providers such as AWS for their Nitro Enclaves (AWS, 2024a) or Apple for their private cloud (Apple, 2024). Virtual TPMs are also the standard for trusted computing in modern clouds (Azure, 2024; Raj et al., 2016; AWS, 2024b). As a result, all these architectures are insecure under our threat model as they all expose sensitive key material to microarchitectural side channels.
@@
Hardware TEE Platforms. To the best of our knowledge, all commercial hardware TEE platforms are vulnerable to microarchitectural side channels. This includes, but is not limited to, Intel SGX (Costan and Devadas, 2016), Intel TDX (Cheng et al., 2024), AND SEV-SNP (SEV-SNP, 2020), ARM TrustZone (Pinto and Santos, 2019) or ARM CCA (Li et al., 2022). Beyond their vulnerability to side channels, commercial platforms also suffer from availability issues as they require dedicated hardware and are not available on consumer-grade processors. For example, Intel announced that SGX would be deprecated on non-server-grade processors (Intel, 2024) and also deprecated its remote attestation service (Corporation, 2024). An exhaustive comparison of existing platforms with Argos can be found in table 7. Academic platform (Costan et al., 2016; Ferraiuolo et al., 2017; Lee et al., 2020b; Bahmani et al., 2021; Feng et al., 2021; Azab et al., 2011; Brasser et al., 2019; Deng et al., 2014; Feng et al., 2021; Van Strydonck et al., 2023; Weiser et al., 2019) are limited in term of how much they can customize the hardware to defend against microarchitectural side channels. As a result, the vast majority simply consider side channels out of scope, with the exception of MI6 (Bourgeat et al., 2019) which is built on top of an open-source RISCV processor.
@@
Combining Cryptography and TEEs. Previous work has identified that protecting remote attestation mechanisms alone against microarchitectural side channels was easier than enforcing confidentiality of entire enclave programs (Tramer et al., 2017). Some even suggest to combine trusted hardware with cryptography to build hybrid mechanisms, but all show important performance overheads, use (insecure) SGX and are vulnerable to microarchitectural side channels. CryptoFlow (Kumar et al., 2020) looks at secure inference and proposes to build maliciously-secure MPC from semi-honest MPC schemes using SGX. They show a 3x performance overhead over the semi-honest scheme. Chex-Mix (Natarajan et al., 2021) leverages SGX with FHE to build secure inference, but considers microarchitectural side channels out of scope and even relies on SGX confidentiality for the privacy of the model weights. They show a 142% performance overhead compared to semi-honest FHE and consider key recovery attacks out of scope. Viand et. al. (Viand et al., 2023) introduces many useful formalism we get inspiration from in Argos. Their paper primarily focuses on combining FHE with cryptographic proofs but they also consider a FHE-in-TEE approach and design a special protocol to optimize its performance. Nevertheless, they use SGX, do not adapt their formalism to the FHE-in-TEE approach, and their results show an 80x performance overhead when compared to Argos. In contrast, Argos presents a new TEE design for malicious and verifiable FHE with a dedicated remote attestation scheme. Argos is secure by desing against microarchitectural side channels and shows minimal performance overhead compared to semi-honest baselines.
@@ 11. Conclusion
We present Argos, the first integrity-only enclave platform designed to build maliciously-secure verifiable FHE. Argos is secure by design against microarchitectural side channels and transient execution attacks and can be used to build fully malicious and authenticated PSI and PIR schemes. It requires no specialized hardware and is compatible with commodity processors from 2008 onward. Argos only incurs 3% average performance overhead for FHE computation, less than 8% performance overhead for complex protocols, and virtually no communication overhead. By demonstrating how to combine cryptography with trusted hardware, Argos paves the way for widespread deployment of FHE-based protocols beyond the semi-honest setting, without the overhead of cryptographic proofs.
@@ Acknowledgements.
The authors thank Charly Castes, Adrien Ghosn, Neelu S. Kalani, and the authors of Tyche (Castes et al., 2023) for providing early access to their code and for their invaluable assistance in setting up Tyche on our hardware and adapting the platform to our needs. We thank Sacha Servan-Schreiber for helpful discussions on fully homomorphic encryption and useful feedback. The authors used Claude to revise the text in most of this paper to correct for typos, grammatical errors, and awkward phrasing.References
@@
- Apple (2024) Apple. 2024. Private Cloud Compute: A new frontier for AI privacy in the cloud. https://security.apple.com/blog/private-cloud-compute/.
@@ Appendix A Primer on Measured Boot
Measure boot is obtained using a chain of trust that ties the measurement of the operating system or hypervisor to a component that is inherently trusted by the system (the root). Starting from the root, every element of the chain measures (hashes) and signs the binary of the next element, ensuring its integrity. This process is sometimes referred to as secure boot, as, historically, the chain of trust was composed of the entire boot chain. The root of trust (RoT) is the base case of this recursive security argument, often a hardware mechanism (e.g., read only memory) that enforces integrity for the first instructions executed at boot and which integrity is supported by a certificate signed by the manufacturer (Ermolov, 2016; Zimmer and Krau, 2016).
@@
When elements of the boot chain are measured (for instance, before any memory can be trusted), these measurements need to be stored in a secure location where it cannot be tempered with, for instance by a compromised BIOS. To serve that functionality, a small chip or coprocessor called the trusted platform module (TPM) will be in charge of securely storing hash-based measurements in special registers (PCRs). The interface exposed by the TPM is quite limited. For example, measurements can only be hash-extended (never reset). TPMs also have other capabilities, like long-term secrets storage referred to a secret sealing (e.g., a key for disk encryption (Tan et al., 2020)) that can only be accessed if the PCRs are in a specific state, i.e., the correct system has been booted. It also has attestation capabilities, which will use a private key tied to the TPM to sign or attest the values of the PCRs and the state of the system. Nevertheless, the TPM is merely a subordinate to the RoT and the software in the chain of trust, as it needs to be instructed and fed inputs to extend its internal measurements.

### v2

@@ 9.7. Analysis of the Remaining Attack Surface
Argos primarily defends against microarchitectural side channels, which constitute the vast majority of TEE attacks so far (43 published according to (Li et al., 2024)). While Argos is vulnerable to fault injection (Qiu et al., 2019; Murdock et al., 2020; Kenjar et al., 2020) and Rowhammer attacks (less studied on TEEs), these threats are not fatal – software can be hardened against fault injections (Breier and Hou, 2022), and no Rowhammer attacks have demonstrated the capabilities needed to compromise Argos (gaining hypervisor privilege on a hardware VM). Regarding physical attacks, Argos protects against ColdBoot (Halderman et al., 2009) and physical side channels on the CPU (Lee et al., 2020a), but remains vulnerable to physical fault injection attacks (Chen et al., 2021; Buhren et al., 2021). The TPM only exposes a minimal attack surface for physical side channels, and modern TPMs are now directly integrated into the SoC, significantly raising the bar for these attacks (Jacob et al., 2023). Indeed, integrated TPMs do not implement modern features such as DVFS or RAPL, making all known software attacks such as Hertzbleed (Wang et al., 2022) or PLATYPUS (Lipp et al., 2021) and others (Zhang et al., 2021) impossible. Other physical attacks are out-of-scope, but would still be significantly less practical: 1) CPU-noise most likely masks signals from small integrated TPM, 2) TPM cryptography is assumed constant-time, thwarting simple power and EM attacks (Mantel et al., 2018; Genkin et al., 2015), 3) for constant-time code, side channels are sometimes due to advanced circuit optimizations (Lipp et al., 2022), unlikely in a small integrated co-processor. Finally, our TCB might still contain bugs, and attacks that exploit implementation mistakes have been demonstrated on the Intel DRoT and the associated secure coprocessor (Wojtczuk and Rutkowska, 2009; Wojtczuk et al., 2009; Ermolov and Goryachy, 2017; Corporation, 2019; Buhren et al., 2019). However, these implementation errors can be patched when discovered —unlike vulnerabilities to microarchitectural side channel attacks.
@@
SC: Side Channel, P: Protected, V: Vulnerable. +’s and -’s represent relative (and subjective) measure of performance.
@@ 10. Related Work
TPMs and Hypervisor-based TEEs. Trusted platform modules (TPMs) combined with dynamic root-of-trust technologies have long served as dedicated hardware to enforce platform integrity. Flicker (McCune et al., 2008) proposes to use these two technologies combined to isolate and attest small pieces of application logic, or PALs while reducing the software TCB to a handful of lines of code. No keys ever leave the TPM in Flicker, which gives it the same resistance to microarchitectural side channels as Argos (even if not discussed in the paper at the time). However, their complete reliance on hardware mechanisms creates significant overheads, i.e., 2-3 orders of magnitude. Virtualization technologies (making their appearance in processors around 2008) aim at isolating different trusted domains in the form of virtual machines (VMs). They create a new privileged execution mode dedicated to the hypervisor and provide hardware primitives such as extended (or nested) page tables to enforce memory isolation between different security domains. These technologies identified the opportunity to virtualize the TPM (Perez et al., 2006) and provide confidentiality and integrity for trusted domains with minimal performance overhead. TrustVisor (McCune et al., 2010) is the first work to introduce this hypervisor-based architecture, which was eventually adopted by cloud providers such as AWS for their Nitro Enclaves (AWS, 2024a). Virtual TPMs are also the standard for trusted computing in modern clouds (Azure, 2024; Raj et al., 2016; AWS, 2024b). As a result, all these architectures are insecure under our threat model as they expose sensitive key material to microarchitectural side channels.
@@
Hardware TEE Platforms. All commercial hardware TEE platforms are vulnerable to microarchitectural side channels. This includes, but is not limited to, Intel SGX (Costan and Devadas, 2016), Intel TDX (Cheng et al., 2024), AMD SEV-SNP (SEV-SNP, 2020), ARM TrustZone (Pinto and Santos, 2019) and ARM CCA (Li et al., 2022). Secrets and key material are sometimes stored on secure co-processors, such as Intel ME (Corporation, 2022), AMD PSP (Buhren et al., 2019), or the Apple secure enclave (Apple, 2022). However, to our knowledge, all existing designs expose some secret material on the CPU such as platform-level secrets for Intel SGX/TDX (Costan and Devadas, 2016), vTPM key material for AMD SEV-SNP (Antonino et al., 2023) or application-level keys for Apple Private Cloud Compute (Apple, 2024b). Beyond their vulnerability to side channels, commercial platforms also suffer from availability issues as they require dedicated hardware and are not available on consumer-grade processors. For example, Intel announced that SGX would be deprecated on non-server-grade processors (Intel, 2024) and also deprecated its remote attestation service (Corporation, 2024). A comprehensive comparison of existing platforms with Argos can be found in Table 7. Academic platforms (Costan et al., 2016; Ferraiuolo et al., 2017; Lee et al., 2020b; Bahmani et al., 2021; Feng et al., 2021; Azab et al., 2011; Brasser et al., 2019; Deng et al., 2014; Feng et al., 2021; Van Strydonck et al., 2023; Weiser et al., 2019) are limited in terms of how much hardware customization they can perform to defend against microarchitectural side channels. As a result, the vast majority simply consider side channels out of scope, with the exception of Sanctum (Costan et al., 2016) and MI6 (Bourgeat et al., 2019) which are built on top of an open-source RISCV processor.
@@
Combining Cryptography and TEEs. Previous work has identified that protecting remote attestation mechanisms alone against microarchitectural side channels was easier than enforcing confidentiality of entire enclave programs (Tramer et al., 2017). Some even suggest to combine trusted hardware with cryptography to build hybrid mechanisms, but all show important performance overheads, use (insecure) SGX and are vulnerable to microarchitectural side channels. CrypTFlow (Kumar et al., 2020) looks at secure inference and proposes to build maliciously-secure MPC from semi-honest MPC schemes using SGX. They show a 3 performance overhead over the semi-honest scheme. Chex-Mix (Natarajan et al., 2021) leverages SGX with FHE to build secure inference, but considers microarchitectural side channels out of scope and even relies on SGX confidentiality for the privacy of the model weights. They show a 142% performance overhead compared to semi-honest FHE and consider key recovery attacks out of scope. Viand et. al. (Viand et al., 2023) introduces many useful formalisms that inspired Argos. Their paper primarily focuses on combining FHE with cryptographic proofs but they also consider an FHE-in-TEE approach and design a special protocol to optimize its performance. Nevertheless, they use SGX, do not adapt their formalism to the FHE-in-TEE approach, and their results show an 80 performance overhead when compared to Argos. In contrast, Argos presents a new TEE design for malicious and verifiable FHE with a dedicated remote attestation scheme. Argos is secure by design against microarchitectural side channels and shows minimal performance overhead compared to semi-honest baselines.
@@ 11. Conclusion
We present Argos, the first integrity-only enclave platform designed to build maliciously-secure verifiable FHE. Argos is secure by design against microarchitectural side channels and transient execution attacks and can be used to build fully malicious and authenticated PSI and PIR schemes. It requires no specialized hardware and is compatible with commodity processors from 2008 onward. Argos only incurs 3% average performance overhead for FHE computation, less than 8% performance overhead for complex protocols, and no communication overhead. By demonstrating how to combine cryptography with trusted hardware, Argos paves the way for widespread deployment of FHE-based protocols beyond the semi-honest setting, without the overhead of cryptographic proofs.
@@ Acknowledgements.
The authors thank Charly Castes, Adrien Ghosn, Neelu S. Kalani, and the authors of Tyche (Castes et al., 2023) for providing early access to their code and for their invaluable assistance in setting up Tyche on our hardware and adapting the platform to our needs. We thank Sacha Servan-Schreiber for helpful discussions on fully homomorphic encryption and useful feedback. The authors used Claude to revise the text in most of this paper to correct for typos, grammatical errors, and awkward phrasing. This research was supported in part by NSF contracts 2330065, 2115587 and 1955270, and Natural Sciences and Engineering Research Council of Canada (NSERC) under funding reference number RGPIN-2023-04796, and an NSERC-CSE Research Communities Grant under funding reference number ALLRP-588144-23. Any research, opinions, or positions expressed in this work are solely those of the authors and do not represent the official views of the Communications Security Establishment Canada or the Government of Canada.References
@@
- Agrawal et al. (2003) Dakshi Agrawal, Bruce Archambeault, Josyula R Rao, and Pankaj Rohatgi. 2003. The EM side—channel (s). In Cryptographic Hardware and Embedded Systems-CHES 2002: 4th International Workshop Redwood Shores, CA, USA, August 13–15, 2002 Revised Papers 4. Springer, 29–45.
- Almeida et al. (2016) José Bacelar Almeida, Manuel Barbosa, Gilles Barthe, François Dupressoir, and Michael Emmi. 2016. Verifying Constant-Time Implementations. In 25th USENIX Security Symposium (USENIX Security 16). 53–70.
@@
- Antonino et al. (2023) Pedro Antonino, Ante Derek, and Wojciech Aleksander Wołoszyn. 2023. Flexible remote attestation of pre-SNP SEV VMs using SGX enclaves. IEEE access 11 (2023), 90839–90856.
@@
- Apple (2024a) Apple. 2024a. Private Cloud Compute: A new frontier for AI privacy in the cloud. https://security.apple.com/blog/private-cloud-compute/.
- Apple (2024b) Apple. 2024b. Private Cloud Compute Security Guide. https://security.apple.com/documentation/private-cloud-compute.
@@
- Backes et al. (2010) Michael Backes, Markus Dürmuth, Sebastian Gerling, Manfred Pinkal, Caroline Sporleder, et al. 2010. Acoustic Side-Channel attacks on printers. In 19th USENIX Security Symposium (USENIX Security 10).
@@
- Brakerski et al. (2014) Zvika Brakerski, Craig Gentry, and Vinod Vaikuntanathan. 2014. (Leveled) fully homomorphic encryption without bootstrapping. ACM Transactions on Computation Theory (TOCT) 6, 3 (2014), 1–36.
@@
- Buhren et al. (2019) Robert Buhren, Christian Werling, and Jean-Pierre Seifert. 2019. Insecure until proven updated: analyzing AMD SEV’s remote attestation. In Proceedings of the 2019 ACM SIGSAC Conference on Computer and Communications Security. 1087–1099.
@@
- Genkin et al. (2015) Daniel Genkin, Itamar Pipman, and Eran Tromer. 2015. Get your hands off my laptop: physical side-channel key-extraction attacks on pcs: Extended version. Journal of Cryptographic Engineering 5 (2015), 95–112.
@@
- Kocher et al. (2020) Paul Kocher, Jann Horn, Anders Fogh, Daniel Genkin, Daniel Gruss, Werner Haas, Mike Hamburg, Moritz Lipp, Stefan Mangard, Thomas Prescher, et al. 2020. Spectre attacks: Exploiting speculative execution. Commun. ACM 63, 7 (2020), 93–101.
- Koruyeh et al. (2018) Esmaeil Mohammadian Koruyeh, Khaled N Khasawneh, Chengyu Song, and Nael B Abu-Ghazaleh. 2018. Spectre Returns! Speculation Attacks using the Return Stack Buffer.. In WOOT@ USENIX Security Symposium.
@@
- Lipp et al. (2022) Moritz Lipp, Daniel Gruss, and Michael Schwarz. 2022. AMD prefetch attacks through power and time. In 31st USENIX Security Symposium (USENIX Security 22). 643–660.
- Lipp et al. (2021) Moritz Lipp, Andreas Kogler, David Oswald, Michael Schwarz, Catherine Easdon, Claudio Canella, and Daniel Gruss. 2021. PLATYPUS: Software-based Power Side-Channel Attacks on x86. In 2021 IEEE Symposium on Security and Privacy (SP). IEEE.
- Liu et al. (2015) Fangfei Liu, Yuval Yarom, Qian Ge, Gernot Heiser, and Ruby B Lee. 2015. Last-level cache side-channel attacks are practical. In 2015 IEEE symposium on security and privacy. IEEE, 605–622.
- Loughlin et al. (2023) Kevin Loughlin, Jonah Rosenblum, Stefan Saroiu, Alec Wolman, Dimitrios Skarlatos, and Baris Kasikci. 2023. Siloz: Leveraging DRAM isolation domains to prevent inter-VM rowhammer. In Proceedings of the 29th Symposium on Operating Systems Principles. 417–433.
- Mantel et al. (2018) Heiko Mantel, Johannes Schickel, Alexandra Weber, and Friedrich Weber. 2018. How secure is green IT? The case of software-based energy side channels. In Computer Security: 23rd European Symposium on Research in Computer Security, ESORICS 2018, Barcelona, Spain, September 3-7, 2018, Proceedings, Part I 23. Springer, 218–239.
@@
- Osvik et al. (2006) Dag Arne Osvik, Adi Shamir, and Eran Tromer. 2006. Cache attacks and countermeasures: the case of AES. In Topics in Cryptology–CT-RSA 2006: The Cryptographers’ Track at the RSA Conference 2006, San Jose, CA, USA, February 13-17, 2005. Proceedings. Springer, 1–20.
@@
- Schwarz et al. (2019) Michael Schwarz, Martin Schwarzl, Moritz Lipp, Jon Masters, and Daniel Gruss. 2019. Netspectre: Read arbitrary memory over network. In Computer Security–ESORICS 2019: 24th European Symposium on Research in Computer Security, Luxembourg, September 23–27, 2019, Proceedings, Part I 24. Springer, 279–299.
@@
- Yarom and Falkner (2014) Yuval Yarom and Katrina Falkner. 2014. FLUSH+ RELOAD: A high resolution, low noise, L3 cache side-channel attack. In 23rd USENIX Security Symposium (USENIX Security 14). 719–732.
- Zhang et al. (2024) Zhi Zhang, Decheng Chen, Jiahao Qi, Yueqiang Cheng, Shijie Jiang, Yiyang Lin, Yansong Gao, Surya Nepal, Yi Zou, Jiliang Zhang, et al. 2024. SoK: Rowhammer on Commodity Operating Systems. In Proceedings of the 19th ACM Asia Conference on Computer and Communications Security. 436–452.
- Zhang et al. (2021) Zhenkai Zhang, Sisheng Liang, Fan Yao, and Xing Gao. 2021. Red alert for power leakage: Exploiting intel rapl-induced side channels. In Proceedings of the 2021 ACM Asia Conference on Computer and Communications Security. 162–175.
@@ Appendix A Primer on Measured Boot
Measured boot is obtained using a trust chain that binds the measurement of the operating system or hypervisor to a component that is inherently trusted by the system (the root). Starting from the root, every element of the chain measures (hashes) and signs the binary of the next element, ensuring its integrity. This process is sometimes referred to as secure boot, as, historically, the chain of trust was composed of the entire boot chain. The root of trust (RoT) is the base case of this recursive security argument, often a hardware mechanism (e.g. read-only memory) that enforces integrity for the first instructions executed at boot and which integrity is supported by a certificate signed by the manufacturer (Ermolov, 2016; Zimmer and Krau, 2016).
@@
When elements of the boot chain are measured (for instance, before any memory can be trusted), these measurements need to be stored in a secure location where they cannot be tampered with, for instance, by a compromised BIOS. To provide that functionality, a small chip or coprocessor, called a trusted platform module (TPM) will be in charge of securely storing hash-based measurements in special registers (PCRs). The interface exposed by the TPM is quite limited. For example, measurements can only be hash-extended (never reset). TPMs also have other capabilities, like long-term secrets storage referred to as secret sealing (e.g., a key for disk encryption (Tan et al., 2020)) that can only be accessed if the PCRs are in a specific state, i.e., the correct system has been booted. The TPM also has attestation capabilities, which will use a private key tied to the TPM to sign or attest the values of the PCRs and the state of the system. However, the TPM is merely a subordinate to the RoT and the software in the chain of trust, as it needs to be instructed and fed inputs to extend its internal measurements.